OpenStudy (anonymous):

the sum of four consecutive odd integers is 56.find the largest of these integers

OpenStudy (nottim):

56=n+2n+4n

OpenStudy (nottim):

dat the formula? i dunno

OpenStudy (anonymous):

x+x+2+x+4+x+6=56 this this and get ur anwer.. cos there is a differece of 2 in each consective odd nmber

OpenStudy (nottim):

oh. well shucks then

OpenStudy (anonymous):

no @NotTim u r wrong .. there is differece of two between each odd nmber so we will take it as x+(x+2)+(x+4)+(x+6)=56 the 4th and largest odd nmber will be x+6 .. n r welkum any time

OpenStudy (nottim):

grrrrr

OpenStudy (anonymous):

ok